# ModHiFi: Identifying High Fidelity predictive components for Model Modification

### v1

@@ 5.3 Effectiveness of HiFi components: Exploring Q2
To answer Q2, and verify whether HiFi components are the components that matter for the final predictive performance, we measure the effects of the fidelity of a component getting destroyed by noising. For a ResNet-50 on CIFAR-10, when 20% of the HiFi components are perturbed with a zero mean Gaussian noise with standard deviation of 0.01, the accuracy of the model drops by around 12%, whereas perturbing 80% of the non-HiFi components identically results in an accuracy drop of only 1%. At 50% of components with a noise of standard deviation 0.02, the accuracy drops by 85% when HiFi components are noise compared to only around 1.4% when non-HiFi components are noised. In Section˜C.1.3, we make similar observations across various models and tasks. In Section˜C.1.2, we additional experiments where we compare the removal of HiFi, non HiFi, and random sets of the same size and make similar observations.
@@ 5.4 Structured Pruning Experiments: Q3
GReg-2 [wang2021neural] 73.9 3.02x 2.31x 1.36x 1.53x
OTO [chen2021trainonceoneshotneural] 74.7 2.86x 2.81x 1.25x 1.45x
DepGRAPH [fang2023depgraphstructuralpruning] 75.83 2.07x - - -
ThiNet [luo2017thinet] 71.6 3.46x 2.95x 1.38x 1.50x
DFPC (30) [narshana2022dfpc] 75.9 1.98x 1.84x 1.42x 1.53x
DFPC (54) [narshana2022dfpc] 73.80 3.46x 2.65x 2.37x 2.38x
@@
DFPC [narshana2022dfpc] 90.25 1.46x 2.07x
$L_{2}$ [li2017pruning] 15.91 4.07x 4.71x
$L_{2}$ w/ ST [li2017pruning] 90.12 4.07x 4.71x
@@
5.4.1 Vision Models
Baselines
We compare against the state of the art structured pruning algorithms specialized for pruning vision models [narshana2022dfpc, luo2017thinet, wang2021neural, chen2021trainonceoneshotneural], and present additional results on other architectures and datasets in Section˜C.4 where we make similar observations. Following [2208.11580], we update the batch norm statistics using the data from distributional access.
@@ 5.4.1 Vision Models
We find that our method results in better accuracy-vs-sparsity trade off when compared to other algorithms across datasets. We also train a model obtained with $L_{2}$ norm-based structured pruning using the synthetic set based on CIFAR10 for comparison. We observe that for the same FLOP sparsity, our method obtains higher accuracy than the model finetuned on synthetic samples, indicating that our method is able to outperform fine-tuning in some cases using synthetic samples for the same sparsity. For the ImageNet dataset, we compare against various state-of-the-art structured pruning algorithms for networks with complex interconnections with training. In the training regime, we observe that for models of similar accuracy, our algorithm obtains the best accuracy-speedup tradeoff with fewer epochs of finetuning. Details of pre-trained networks and post-training are given in Section˜C.7.2. Our study of the effect of the quality of synthetic samples on our algorithm in Section˜C.3.3 indicates that the sparsity-accuracy tradeoff of our algorithm degrades with lower quality samples, but it does not degrade as much as $L_{2}$ pruning + finetuning on synthetic samples.
@@ 5.4.2 Large Language Models
20% ShortGPT [men2024shortgpt] 14.32 58.33 38.05 72.58 65.51 65.27 59.95
SliceGPT [ashkboos2024slicegpt] 8.13 50.08 31.14 64.85 62.04 48.84 51.39
@@
30% ShortGPT [men2024shortgpt] 33.21 48.65 32.85 64.31 64.33 56.13 53.25
SliceGPT [ashkboos2024slicegpt] 10.96 44.19 27.47 58.71 57.46 41.27 45.82
@@
Baselines
We evaluate ModHiFi on Llama-2-7B, comparing it against state-of-the-art algorithms for structured pruning [ashkboos2024slicegpt, men2024shortgpt]. The use of calibration datasets to compute statistics is in line with our framing of distributional access to data, since LLMs do not make their training data openly accessible. Unless otherwise specified, the algorithms use WikiText-2 for calibration, with 128 samples of length 1024  [^3] ShortGPT’s calibration data is not publicly available.. None of the algorithms perform post-pruning recovery fine-tuning. Additional details about our choice of baselines can be found in Section˜C.4.3.
@@
Swin-T [liu2021swintransformerhierarchicalvision] Base 92.31 92.31 -
Jia et al. [jia2023model] 1.20 90.69 235
@@
Evaluation
We also measure the performance of the model via its zero-shot accuracy on a suite of standard NLP tasks [clark2018thinksolvedquestionanswering, sakaguchi2019winograndeadversarialwinogradschema, zellers2019hellaswagmachinereallyfinish, bisk2019piqareasoningphysicalcommonsense] and WikiText perplexity. In Table˜3, we see that our method is competitive, with consistently high average and task-specific performance, and outperforming at moderate sparsities. We find that the quality of the calibration set plays an important role, with the performance of ModHiFi-P-Alpaca being consistently higher than that of ModHiFi-P-WikiText. This tells us that retaining only HiFi components offers a model-agnostic approach to structured pruning, with its application to LLMs requiring no changes from its application to vision models.
5.5 Class Unlearning Experiments: Q4
Baselines and Metrics
We report the forget and retain accuracy averaged across 10 classes of the CIFAR10 dataset on ResNet-50 and Swin-T models. We benchmark against Gradient Ascent and jia2023model which are both retraining-based technique for Unlearning.
Unlearning Results
We report the results of our algorithm in Table˜4. To answer Q4, we observe that it is possible to perform unlearning without finetuning in a general editing framework $10\times$ faster than our baseline. In Section˜C.5, we compare results with finetuning using synthetic and training data. We note that the results for Swin-Transformer without finetuning fail to achieve state of the art. However, as reported in Section˜C.5, we observe a drastic improvement with only 3 epochs of finetuning on synthetic samples. With 10 epochs of finetuning with our algorithm, we find that compared to [jia2023model] (who use full training), our forget accuracy is superior when using synthetic samples, and both forget and remain accuracies are superior when using training samples. Experiments with VGG-19 are present in Section˜C.5 where we make similar observations.
@@ 6 Discussion and Conclusion
Limitations and Future Work Our theoretical bounds in Section˜3.2 rely on the local Lipschitz continuity of the network. While we demonstrate that this property holds for well-trained models (including Transformers on bounded domains), it is not guaranteed at initialization. This suggests that the emergence of High-Fidelity components is a consequence of the training dynamic itself. In this work, we use the expected square loss as a measure of distributional similarity, and we leave for future work the exploration of other metrics of distributional similarities like the TV distance or Wasserstein metric.
@@
We authors gratefully acknowledge AMD for their support. The authors thank Ramaswamy Govindarajan (IISc) and Prakash Raghavendra (AMD) for their insight and assistance in this work. The authors are also grateful to the reviewers of this work for improve its content.
@@ 13. Answer: [Yes]
  14. Justification: Yes, for each theoretical result, we provide a complete set of assumptions and a correct proof. Proofs are attached in the appendix. We link the relevant appendices in the main body for the reader.
@@ 18. Answer: [Yes]
  19. Justification: Yes, we fully disclose all the information to reproduce the main experimental results in Section 5 and the appendices mentioned within that section. Moreover, we provide our code, and the data sets used are open-sourced.
@@ 23. Answer: [Yes]
  24. Justification: The code and the instructions to reproduce the experiments are provided in our Anonymous GitHub link provided in the paper.
@@ 28. Answer: [Yes]
  29. Justification: All details to understand the results and reproduce the results are provided in Section 5 and the appendices mentioned in this section.
@@ 38. Answer: [Yes]
  39. Justification: We sufficiently describe the compute resources used for our experiments in Section 5 and the appendices referred to within the section.
@@ 48. Answer: [No]
  49. Justification: We do not discuss the societal impact of the work performed in this manuscript since this is foundational research and not tied to particular applications.
@@ 53. Answer: [N/A]
  54. Justification: We do not release models or data in this work.
@@ 80. Guidelines:
      • In Section˜C.4.2 we compare the pruned ImageNet models of ModHiFi-P against SoTA data-free structured pruning DFPC [narshana2022dfpc] to compare layer-wise sparsity to see where is improved speedup coming from.
@@ Appendix A Related Work
Structured pruning has been widely researched, with a wide variety of methods proposed for it [hoefler2021sparsity]. Unlike unstructured pruning, which sparsifies the weight matrices without changing the architecture of the model [lecun1989optimal, 2208.11580, blalock2020state, tanaka2020pruning, frankle2018lottery], structured pruning enables immediate improvements in real-world performance measures such as inference time and memory footprint without requiring specialized hardware or software [hoefler2021sparsity, molchanov2016pruning, prakash2019optimizing]. A variety of methods have been proposed for structured pruning of convolutional networks, including using norms of weight tensors [li2016pruning, li2017pruning], directional derivative scores [molchanov2016pruning, molchanov2019importance, shen2022structural], feature map ranks [lin2020hrank, sui2021chip], coresets [liebenwein2019provable, baykal2018data, tukan2022pruning], discriminative ability of filters [murti2022tvsprune, liu2021discrimination], and reconstruction error [yu2018nisp, elhalabi2022dataefficient, shuan2025numerical]. However, modern neural networks possess complex interconnections, making them difficult to structurally compress [liu2021group, narshana2022dfpc, fang2023depgraph], for which some recent algorithms have been proposed that use gradient information [liu2021group] or bounds on the reconstruction error [yu2018nisp, narshana2022dfpc]. Moreover, pruning without access to either the training data or the loss function is an increasingly important area of research, for which some works have been proposed that use the discriminative ability of filters as a saliency [liu2021discrimination, murti2022tvsprune]. However, none of these works address the problem of pruning large language models.
Pruning of Large Language Models (LLMs) has garnered significant interest in recent years [zhu2024survey]. A variety of unstructured pruning methods have been proposed, such as [frantar2023sparsegpt, sun2023simple]. However, these methods do not provide direct improvements on inference time and memory footprint. Thus, the problem of pruning models with structural interconnections has naturally been applied to pruning LLMs as well, in works such as [ashkboos2024slicegpt, ma2023llm, xia2023sheared, men2024shortgpt]. A key drawback of these works is that most are not applicable to CNNs or other kinds of models. Our work proposes a unified framework for both pruning models with complex interconnections, including transformers and ResNets, as well as classwise unlearning.
@@
Machine unlearning has gained significant interest in recent years, both for data privacy concerns as well as connections to continual learning [bourtoule2021machine, nguyen2022survey, wang2024comprehensive, hooker2019compressed]. Machine unlearning is typically categorized into exact and approximate unlearning [xu2024machine]. Exact unlearning involves training models from scratch without the forget data (the data to be forgotten), or by training modules or experts on subsets of data [yan2022arcane, xiong2023exact]. Approximate unlearning, on the other hand, refers to techniques that approximate exact unlearning via various approaches [izzo2021approximate, xu2024machine]. Machine unlearning can be further classified into sample unlearning (wherein individual samples or random subsets of samples are unlearned)[sekhari2021remember, ullah2023adaptive] or classwise unlearning (where classes or concepts are unlearned) [graves2021amnesiac, gandikota2023erasing]. In this work, we focus on classwise unlearning.
A variety of approaches have been proposed for classwise unlearning [graves2021amnesiac]. Popular methods include fine-tuning the model without data from the forget class [warnecke2021machine, golatkar2020eternal], gradient ascent on the forget set [graves2021amnesiac, thudi2022unrolling], distillation-based approaches [kurmanji2024towards], and influence function based methods [izzo2021approximate]. More recent work studies using sparsity for machine unlearning, such as [jia2023model], which first sparsifies the model, and then applies a fine-tuning-based unlearning algorithm, or [murti2024discedit, wang2022federated], which identify class-discriminative filters in CNNs, and removes them for unlearning. Two key drawbacks of prior art, however, are: first they exclusively address classwise unlearning, and do not address wider problems of model modification. Second, all prior art assumes access to the original training data. Our proposed approach for classwise unlearning differs from prior art because it only requires synthetic class data, uses a variety of granularities for sparsity in unlearning, and is part of a unified approach to model modification.
@@ Appendix B Proofs
We now provide a proof of Section˜3.2. We first state the properties of normalization layer and provide empirical evidence to justify their validity, followed by a restatement of the theorem, and its proof.
@@
$\mathrm{Norm}(\bm{\phi}(\bm{x}))=\bm{\gamma}^{l}\odot\frac{\mathbf{z}}{\|\mathbf{z}\|_{2}}+\bm{\beta}^{l},\quad\text{where }\mathbf{z}=\mathbf{M}\bm{\phi}(\bm{x}).$
@@
A function $f:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is Lipschitz continuous in it’s domain if there exists a positive scalar constant L such that
@@
Let $\mathcal{X}_{r}=\{\mathbf{x}\in\mathbb{R}^{d}\mid\|\mathbf{x}\|_{2}\geq r>0\}$. Define the map $f:\mathcal{X}_{r}\to\mathbb{S}^{d-1}$ as $f(\mathbf{x})=\frac{\mathbf{x}}{\|\mathbf{x}\|_{2}}$. Then $f$ is Lipschitz continuous on $\mathcal{X}_{r}$ with constant $L_{f}=1/r$.
@@
$\|f_{r}(\bm{x})-f_{r}(\bm{y})\|_{2}^{2}$ $=\left\|\frac{\bm{x}}{\|\bm{x}\|}-\frac{\bm{y}}{\|\bm{y}\|}\right\|^{2}_{2}$
@@
$\|\mathbf{x}-\mathbf{y}\|^{2}_{2}$ $\geq\|\mathbf{x}\|\|\mathbf{y}\|\left(2-2\frac{\mathbf{x}^{\top}\mathbf{y}}{\|\mathbf{x}\|\|\mathbf{y}\|}\right)=\|\mathbf{x}\|\|\mathbf{y}\|\,\|g(\mathbf{x})-g(\mathbf{y})\|_{2}^{2}\geq r^{2}\|g(\mathbf{x})-g(\mathbf{y})\|_{2}^{2}.$
Rearranging terms,
$\|f_{r}(\bm{x})-f_{r}(\bm{y})\|_{2}\leq\frac{1}{r}\|\bm{x}-\bm{y}\|$
@@
Let the input to the $l$-th normalization layer satisfy $\|\mathbf{M}\bm{\Phi}(\bm{x})\|_{2}\geq r>0$ almost surely over the data distribution $\mathcal{D}$. Then, normalization layer is Lipschitz continuous over $\mathcal{D}$ with constant:
$L_{\mathrm{norm}}^{l}=\frac{\|\bm{\gamma}^{l}\|_{\infty}}{r}\|\mathbf{M}\|_{2}.$
@@
Let $\mathbf{u}(\mathbf{x})=\mathbf{M}\mathbf{x}$. Then $\|\mathrm{Norm}(\mathbf{x})-\mathrm{Norm}(\mathbf{y})\|_{2}=\|\bm{\gamma}^{l}\odot(g(\mathbf{u}(\mathbf{x}))-g(\mathbf{u}(\mathbf{y})))\|_{2}\leq\|\bm{\gamma}^{l}\|_{\infty}\frac{1}{r}\|\mathbf{M}(\mathbf{x}-\mathbf{y})\|_{2}$. We apply Lemma˜B.3 to complete the proof. ∎
While the lower bound assumption $\|\mathbf{z}\|\geq r$ is technically not guaranteed for all $\mathbf{x}\in\mathbb{R}^{d}$, we empirically verify that for trained networks, activation norms are strictly bounded away from zero. This validates the local Lipschitz property in the region of interest. We show the layer-wise minimum norm of the pre-LayerNorm representations in Figure˜3, estimated on 100 samples from the Alpaca dataset, where for various models we observe the lower bound to be between 0.2 to 60. For clarity of exposition, we only show the layers with the largest and smallest values along with 5 randomly selected layers. Code for generating these plots can be found in Appendix˜C. We also observe that this value tends to increase for layers deeper in the network, and leave the utilization of this observation to future work.

### v2

@@ 5.3 Effectiveness of HiFi components: Exploring Q2
To answer Q2, and verify whether HiFi components are the components that matter for the final predictive performance, we measure the effects of the fidelity of a component getting destroyed by noising. For a ResNet-50 on CIFAR-10, when 20% of the HiFi components are perturbed with a zero mean Gaussian noise with standard deviation of 0.01, the accuracy of the model drops by around 12%. In contrast, perturbing 80% of the non-HiFi components identically results in an accuracy drop of only 1%. At 50% of components with a noise of standard deviation 0.02, the accuracy drops by 85% when HiFi components are noised compared to only around 1.4% when non-HiFi components are noised. In Section˜C.1.3, we make similar observations across various models and tasks. In Section˜C.1.2, we additionally performed experiments where we compare the removal of HiFi, non HiFi, and random sets of the same size and make similar observations.
@@ 5.4 Structured Pruning Experiments: Q3
5.4.1 Vision Models
Baselines
We compare against the state-of-the-art structured pruning algorithms specialized for pruning vision models [58, 47, 83, 8], and present additional results on other architectures and datasets in Section˜C.4 where we make similar observations. Following [17], we update the batch norm statistics using the data from distributional access.
@@ 5.4.1 Vision Models
GReg-2 [83] 73.9 3.02x 2.31x 1.36x 1.53x
OTO [8] 74.7 2.86x 2.81x 1.25x 1.45x
DepGRAPH [14] 75.83 2.07x - - -
ThiNet [47] 71.6 3.46x 2.95x 1.38x 1.50x
DFPC (30) [58] 75.9 1.98x 1.84x 1.42x 1.53x
DFPC (54) [58] 73.80 3.46x 2.65x 2.37x 2.38x
@@
DFPC [58] 90.25 1.46x 2.07x
$L_{2}$ [41] 15.91 4.07x 4.71x
$L_{2}$ w/ ST [41] 90.12 4.07x 4.71x
@@
We find that our method yields a better accuracy-vs-sparsity tradeoff compared to other algorithms across various datasets. We also train a model obtained with $L_{2}$ norm-based structured pruning using the synthetic set based on CIFAR10 for comparison. In Table˜2, we observe that for the same FLOP sparsity, our method obtains higher accuracy than the model finetuned on synthetic samples, indicating that our method can outperform finetuning in some cases using synthetic samples for the same sparsity. For the ImageNet dataset, we compare our approach against various state-of-the-art structured pruning algorithms for networks with complex interconnections, including those trained on the ImageNet training set. In Section˜5.4.1, we observe that for models of similar accuracy, our algorithm obtains the best accuracy-speedup tradeoff with fewer epochs of finetuning. Details of pre-trained networks and post-training are given in Section˜C.7.2. Our study of the effect of the quality of synthetic samples on our algorithm in Section˜C.3.3 indicates that the sparsity-accuracy tradeoff of our algorithm degrades with lower quality samples, but it does not degrade as much as $L_{2}$ pruning + finetuning on synthetic samples.
@@ 5.4.2 Large Language Models
Baselines We evaluate ModHiFi on Llama-2-7B, comparing it against state-of-the-art algorithms for structured pruning [2, 49]. The use of calibration datasets to compute statistics aligns with our framing of distributional access to data, as LLMs do not make their training data openly accessible. Unless otherwise specified, the algorithms use WikiText-2 for calibration, with 128 samples of length 1024  [^3] ShortGPT’s calibration data is not publicly available.. None of the algorithms performs post-pruning recovery finetuning. Additional details about our choice of baselines can be found in Section˜C.4.3.
Evaluation We also measure the performance of the model via its zero-shot accuracy on a suite of standard NLP tasks [9, 65, 96, 5] and WikiText perplexity. In Table˜3, we observe that our method is competitive, with consistently high average and task-specific performance, and outperforms at moderate sparsity levels. We find that the quality of the calibration set plays a crucial role, with the performance of ModHiFi-P-Alpaca outperforming that of ModHiFi-P-WikiText. This indicates that retaining only HiFi components provides a model-agnostic approach to structured pruning, with its application to LLMs requiring no modifications beyond its application to vision models.
5.5 Class Unlearning Experiments: Q4
Baselines and Metrics
We report the forget and retain accuracy averaged across 10 classes of the CIFAR10 dataset on ResNet-50 and Swin-T models. We benchmark against Gradient Ascent and Jia et al. [33], which are both retraining-based techniques for Unlearning.
Unlearning Results We report the results of our algorithm in Table˜4. To answer Q4, we observe that it is possible to perform unlearning without finetuning in a general editing framework $10\times$ faster than our baseline. In Section˜C.5, we compare results with finetuning using synthetic and training data. We note that the results for Swin-Transformer without finetuning fail to achieve the state of the art. However, as reported in Section˜C.5, we observe a drastic improvement with only three epochs of finetuning on synthetic samples. After 10 epochs of finetuning with our algorithm, we find that our forget accuracy is superior to that of [33] (who use full training) when using synthetic samples. Both forget and remain accuracy are superior when using training samples. Experiments with VGG-19 are present in Section˜C.5 where we make similar observations.
@@ 5.5 Class Unlearning Experiments: Q4
20% ShortGPT [49] 14.32 58.33 38.05 72.58 65.51 65.27 59.95
SliceGPT [2] 8.13 50.08 31.14 64.85 62.04 48.84 51.39
@@
30% ShortGPT [49] 33.21 48.65 32.85 64.31 64.33 56.13 53.25
SliceGPT [2] 10.96 44.19 27.47 58.71 57.46 41.27 45.82
@@
Swin-T [46] Base 92.31 92.31 -
Jia et al. [33] 1.20 90.69 235
@@ 6 Discussion and Conclusion
Limitations and Future Work Our theoretical bounds in Section˜3.2 rely on the local Lipschitz continuity of the network. While we demonstrate that this property holds for well-trained models (including Transformers on bounded domains), it is not guaranteed at initialization. This suggests that the emergence of High-Fidelity components is a consequence of the training dynamics. In this work, we use the expected square loss as a measure of distributional similarity, and we leave for future work the exploration of other metrics of distributional similarity, like the TV distance or Wasserstein metric.
@@
We, the authors, gratefully acknowledge AMD for its support. The authors thank Ramaswamy Govindarajan (IISc) and Prakash Raghavendra (AMD) for their insight and assistance in this work. The authors are also grateful to the reviewers of this work for their valuable feedback, which has significantly improved the content.
@@ 13. Answer: [Yes]
  14. Justification: Yes, for each theoretical result, we provide a complete set of assumptions and a correct proof. Proofs are attached in the appendix, Appendix˜B. We link the relevant appendices in the main body for the reader.
@@ 18. Answer: [Yes]
  19. Justification: Yes, we fully disclose all the information to reproduce the main experimental results in Section 5 and the appendices mentioned within that section.
@@ 23. Answer: [Yes]
  24. Justification: The code and the instructions to reproduce the experiments are provided at the GitHub: https://github.com/DhruvaKashyap/modhifi. Moreover, the data sets used are open-sourced, and details on how to obtain them are provided on our GitHub.
@@ 28. Answer: [Yes]
  29. Justification: All details to understand the results and reproduce the results are provided in Section 5 and the appendices mentioned therein.
@@ 38. Answer: [Yes]
  39. Justification: We sufficiently describe the compute resources used for our experiments in Section 5 and the appendices referred to within the section, specifically, Section˜C.6.
@@ 48. Answer: [No]
  49. Justification: We do not discuss the societal impact of the work performed in this manuscript since this is foundational research and not tied to any particular applications.
@@ 53. Answer: [N/A]
  54. Justification: We do not release generative models or data in this work.
@@ 80. Guidelines:
      • In Section˜C.4.2 we compare the pruned ImageNet models of ModHiFi-P against SoTA data-free structured pruning DFPC [58] to compare layer-wise sparsity to see where is improved speedup coming from.
@@ Appendix A Related Work
Structured pruning has been widely researched, with a wide variety of methods proposed for it [28]. Unlike unstructured pruning, which sparsifies the weight matrices without changing the architecture of the model [39, 17, 6, 77, 15], structured pruning enables immediate improvements in real-world performance measures such as inference time and memory footprint without requiring specialized hardware or software [28, 53, 62]. A variety of methods have been proposed for structured pruning of convolutional networks, including using norms of weight tensors [40, 41], directional derivative scores [53, 54, 69], feature map ranks [43, 74], coresets [42, 4, 81], discriminative ability of filters [56, 44], and reconstruction error [94, 25, 71]. However, modern neural networks possess complex interconnections, making them difficult to structurally compress [45, 58, 13], for which some recent algorithms have been proposed that use gradient information [45] or bounds on the reconstruction error [94, 58]. Moreover, pruning without access to either the training data or the loss function is an increasingly important area of research, for which some works have been proposed that use the discriminative ability of filters as a saliency [44, 56]. However, none of these works address the problem of pruning large language models.
Pruning of Large Language Models (LLMs) has garnered significant interest in recent years [98]. A variety of unstructured pruning methods have been proposed, such as [16, 75]. However, these methods do not provide direct improvements on inference time and memory footprint. Thus, the problem of pruning models with structural interconnections has naturally been applied to pruning LLMs as well, in works such as [2, 48, 88, 49]. A key drawback of these works is that most are not applicable to CNNs or other kinds of models. Our work proposes a unified framework for both pruning models with complex interconnections, including transformers and ResNets, as well as classwise unlearning.
@@
Machine unlearning has gained significant interest in recent years, both for data privacy concerns as well as connections to continual learning [7, 59, 85, 29]. Machine unlearning is typically categorized into exact and approximate unlearning [90]. Exact unlearning involves training models from scratch without the forget data (the data to be forgotten), or by training modules or experts on subsets of data [91, 89]. Approximate unlearning, on the other hand, refers to techniques that approximate exact unlearning via various approaches [31, 90]. Machine unlearning can be further classified into sample unlearning (wherein individual samples or random subsets of samples are unlearned)[67, 82] or classwise unlearning (where classes or concepts are unlearned) [24, 18]. In this work, we focus on classwise unlearning.
A variety of approaches have been proposed for classwise unlearning [24]. Popular methods include fine-tuning the model without data from the forget class [86, 20], gradient ascent on the forget set [24, 79], distillation-based approaches [38], and influence function based methods [31]. More recent work studies using sparsity for machine unlearning, such as [33], which first sparsifies the model, and then applies a fine-tuning-based unlearning algorithm, or [55, 84], which identify class-discriminative filters in CNNs, and removes them for unlearning. Two key drawbacks of prior art, however, are: first they exclusively address classwise unlearning, and do not address wider problems of model modification. Second, all prior art assumes access to the original training data. Our proposed approach for classwise unlearning differs from prior art because it only requires synthetic class data, uses a variety of granularities for sparsity in unlearning, and is part of a unified approach to model modification.
@@ Appendix B Proofs
We now provide a proof of Section˜3.2. We first state the properties of the normalization layer and provide empirical evidence to justify their validity, followed by a restatement of the theorem and its proof.
@@
$\mathrm{NM}(\bm{\phi}(\bm{x}))=\bm{\gamma}^{l}\odot\frac{\mathbf{z}}{\|\mathbf{z}\|_{2}}+\bm{\beta}^{l},\quad\text{where }\mathbf{z}=\mathbf{M}\bm{\phi}(\bm{x}).$
@@
A function $f:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is Lipschitz continuous in its domain if there exists a positive scalar constant L such that
@@
Let $\mathcal{X}_{r}=\{\mathbf{x}\in\mathbb{R}^{d}\mid\|\mathbf{x}\|_{2}\geq r>0\}$. Define the map $f:\mathcal{X}_{r}\to\mathbb{S}^{d-1}$ as $f(\mathbf{x})=\frac{\mathbf{x}}{\|\mathbf{x}\|_{2}}$. Then $f$ is Lipschitz continuous on $\mathcal{X}_{r}$ with constant $L_{f}=1/r$. That is,
$\|f(\bm{x})-f(\bm{y})\|_{2}\leq\frac{1}{r}\|\bm{x}-\bm{y}\|$
@@
$\|f(\bm{x})-f(\bm{y})\|_{2}^{2}$ $=\left\|\frac{\bm{x}}{\|\bm{x}\|}-\frac{\bm{y}}{\|\bm{y}\|}\right\|^{2}_{2}$
@@
$\|\mathbf{x}-\mathbf{y}\|^{2}_{2}$ $\geq\|\mathbf{x}\|\|\mathbf{y}\|\left(2-2\frac{\mathbf{x}^{\top}\mathbf{y}}{\|\mathbf{x}\|\|\mathbf{y}\|}\right)=\|\mathbf{x}\|\|\mathbf{y}\|\,\|f(\mathbf{x})-f(\mathbf{y})\|_{2}^{2}\geq r^{2}\|f(\mathbf{x})-f(\mathbf{y})\|_{2}^{2}.$
Rearranging terms completes the proof. ∎
Using Lemma˜B.3, we can show that the operation performed by normalization layers is Lipschitz continuous in Corollary˜B.4.
@@
Let the input to the $l$-th normalization layer satisfy $\|\mathbf{M}\bm{\Phi}(\bm{x})\|_{2}\geq r>0$ for all $\bm{x}$. Then, the normalization layer satisfies
$\|\mathrm{NM}(\bm{\phi}(\bm{x}))-\mathrm{NM}(\bm{\phi}(\bm{y}))\|\leq\frac{\|\bm{\gamma}^{l}\|_{\infty}}{r}\|\mathbf{M}\|_{2}\|\bm{\phi}(\bm{x})-\bm{\phi}(\bm{y})\|.$
@@
Let $\mathbf{u}(\mathbf{x})=\mathbf{M}\mathbf{x}$. Then $\|\mathrm{NM}(\mathbf{x})-\mathrm{NM}(\mathbf{y})\|_{2}=\|\bm{\gamma}^{l}\odot\left(\frac{\mathbf{u}(\mathbf{x})}{\|\mathbf{u}(\bm{x})\|}-\frac{\mathbf{u}(\mathbf{y})}{\|\mathbf{u}(\bm{y})\|}\right)\|_{2}\leq\|\bm{\gamma}^{l}\|_{\infty}\frac{1}{r}\|\mathbf{M}(\mathbf{x}-\mathbf{y})\|_{2}$ by applying Lemma˜B.3 to complete the proof. ∎
While the lower bound assumption $\|\mathbf{z}\|\geq r$ is technically not guaranteed for all $\mathbf{x}\in\mathbb{R}^{d}$, we empirically verify that for trained networks, activation norms are strictly bounded away from zero. This validates the local Lipschitz property in the region of interest. We show the layer-wise minimum norm of the pre-LayerNorm representations in Fig.˜3, estimated on 100 samples from the Alpaca dataset. For various models, we observe the lower bound to be between 0.2 and 60. For clarity of exposition, we only show the layers with the largest and smallest values, along with 5 randomly selected layers. Code for generating these plots can be found in Appendix˜C. We also observe that this value tends to increase for layers deeper in the network, and leave the utilization of this observation to future work.